OpenStudy (anonymous):

Given the function f(x) = 2x2 + 15, evaluate f(3)

jimthompson5910 (jim_thompson5910):

replace all copies of x with 3 then evaluate and simplify

jimthompson5910 (jim_thompson5910):

f(x) = 2x^2 + 15 f(3) = 2(3)^2 + 15 f(3) = ???

OpenStudy (anonymous):

51?

OpenStudy (anonymous):

that isnt one of my choices

jimthompson5910 (jim_thompson5910):

this is the wrong way to do it 2*3^2 + 15 6^2 + 15 36 + 15 51

jimthompson5910 (jim_thompson5910):

this is the right way to do it f(3) = 2(3)^2 + 15 f(3) = 2(9) + 15 ... you exponentiate or square before you multiply f(3) = 18 + 15 f(3) = 33

OpenStudy (anonymous):

alright thank you!

jimthompson5910 (jim_thompson5910):

you're welcome, practice practice practice with PEMDAS a lot more because it will show up a lot in algebra